If f(x) is an odd function and f(-3)=6,what is the value of f(6)? can you explain it for me thanks.
Found 2 solutions by venugopalramana, Nate:
Answer by venugopalramana(3286)

If f(x) is an odd function and f(-3)=6,what is the value of f(6)? can you explain it for me thanks.
A FUNCTION IS SAID TO BE ODD IF F(X)=-F(-X)
WE ARE GIVEN F(-3)=6
HENCE F(-(-3))=-6
F(3)=-6...THAT IS ALL..THERE IS NO WAY WE CAN PREDICT F(6)
FOR EX.1......F(X)=-2X IS AN ODD FUNCTION
F(-3)=-2*-3=6...AND.....F(3)=-2*3=-6....SO F(-3)=-F(3)..AND...F(6)=-2*6=-12
EX.2..........F(X)=-2X^3/9...IS AN ODD FUNCTION..F(-3)=-2*-27/9=6
F(3)=-2*27/9=-6...SO F(-3)=-F(3)...AND.......F(6)=-2*216/9=-48...SO F(6) CANNOT BE PREDICTED FROM F(-3)..OR THE FUNCTION BEING ODD

Answer by Nate(3500)

Odd: -f(-x)
-f(-3x) = 6
x = 2
odd: f(x) = -2x
f(6) = -2*6
f(6) = -12
Also:
odd: f(x) = x - 9
f(6) = 6 - 9
f(6) = -3
There are many possibilities.